OpenStudy (anonymous):

what is a counterexample for the conjecture? conjecture:Any number that is divisible by 2 is also divisible by 4. A.32 B.12 C.40 D.18

OpenStudy (anonymous):

can u help me

OpenStudy (irishboy123):

the counter example involves finding a number that can be divided* by 2 but not 4. work through each number and you will find one. *everything is divisible by everything. the missing language here is that the "answer" must be a whole number.

OpenStudy (anonymous):

is it 12

OpenStudy (anonymous):

letter b

jimthompson5910 (jim_thompson5910):

is 12 divisible by 4?

OpenStudy (anonymous):

wait oh so then how do u no which could be divisibl by 4?

jimthompson5910 (jim_thompson5910):

you're looking for a number that's divisible by 2, but not 4 you want to disprove the claim that "Any number that is divisible by 2 is also divisible by 4"

OpenStudy (irishboy123):

dude 12 = 4*3 = 2 *6 look for a number that 2 can go into but not 4.

OpenStudy (anonymous):

ohh 18

OpenStudy (anonymous):

am i right

jimthompson5910 (jim_thompson5910):

correct, 18 is divisible by 2 (ie 2 is a factor of 18) but not divisible by 4 (ie 4 is not a factor of 18)

OpenStudy (anonymous):

thank u very much

jimthompson5910 (jim_thompson5910):

np

OpenStudy (irishboy123):

cool!